A cart of mass M traveling to the right on a frictionless track with a speed 4v0 collides with another cart of mass 2M traveling to the left with a speed v0. If this collision is perfectly elastic, and carts travel in opposite directions after collision, determine the speeds of the two carts immediately after the collision in terms of M and v0
Answer #1

speed v1 and v2 immediatly after collision is

v1=-8Vo/3 (towards left)

v2=7Vo/3 (towards right)

<<<<<<<<<< explanation >>>>>>>>>>>>>>

4,- mi m2 2 Before collision After collision

applying conservation of linear momentum

m1u1+m2u2=m1v1+m2v2

==>m1(u1-v1)=m2(u2-v2) ................................ 1

applying conservation of kinetic energy

1/2m1u12+1/2m2u22=1/2m1v12+1/2m2v22

==>m1(u12-v12)=m2(u22-v22) .............................. 2

dividing equation 2 by 1 we get

u1+v1=u2+v2

==>v2 = u1-u2+v1 ................................. 3

putting equation 3 in 1 we get

m1(u1-v1) = m2(u1-u2+v1-u2)

==>(m1-m2)*u1+2m2u2=(m1+m2)v1

==>v1= (m1-m2)*u1+2m2u2/(m1+m2) ....................4

putting value of v1 in 3 we get

v2 = ((m2-m1)u2+2m1u1)/(m1+m2) ....................5

now given

u1=4Vo m1=M

u2=Vo m2=2M

putting values in above equation we get

v1=-8Vo/3 (towards left)

and

v2=7Vo/3 (towards right)

<<<<<<<<<<<<<<<>>>>>>>>>>>>>>>>>.